The solution of which of the following equations is neither a fraction nor an integer? (a) 2x + 6 = 0 (b) 3x – 5 = 0 (c) 5x – 8 = x + 4 (d) 4x + 7 = x + 2

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

Option (d) is correct here
Consider the equation 4x + 7 = x + 2
Transpose 7 to right hand side then it becomes–7 and x to Left hand side then it becomes - x.

4x–x = 2–7
3x = -5
x = -5/3


Thanks